OpenStudy (anonymous):

find the (a) directrix, (b)the focus and roots of (c)parabola y=x2-5x+4

OpenStudy (anonymous):

directrix=y=-5/2 focus=(5/2,-2) roots=x=4,1 ill attach the graph in a bit.

OpenStudy (anonymous):